OpenStudy (anonymous):

x^4-7x^3-12x^2+176x-320 I have to give the exact values list multiple zeros as necessary so i just need to know the probably 4 zeros of this prob

myininaya (myininaya):

x=4 is a zero

myininaya (myininaya):

so x^4-7x^3-12x^2+176x-320=(x-4)(x^3-3x-24x+80)=0

OpenStudy (anonymous):

there is more then 1 zero isnt there

myininaya (myininaya):

(x-4)(x^3-3x^2-24x+80)=0*

myininaya (myininaya):

yes there are 3 more

OpenStudy (anonymous):

nice!

myininaya (myininaya):

x=4 is again another zero

myininaya (myininaya):

(x-4)(x-4)(x^2+x-20)=0

myininaya (myininaya):

now we have 2 more to find

OpenStudy (anonymous):

+5 - 4

myininaya (myininaya):

-5,4

OpenStudy (anonymous):

oo ya i didnt factor them to 0

myininaya (myininaya):

because x^2+x-20=(x+5)(x-4)

OpenStudy (anonymous):

thank you !

myininaya (myininaya):

so 4 multiplicity 3 and 5

myininaya (myininaya):

oops -5

myininaya (myininaya):

np